OpenStudy (anonymous):

MEDAL & FAN for correct answer!!! If f(x) varies directly with x2, and f(x) = 96 when x = 4, find the value of f(2). 192 72 32 24

OpenStudy (anonymous):

24

OpenStudy (anonymous):

can you explain to me how you got that answer?

OpenStudy (anonymous):

Okay first, we'll find C For x=4 we have f(x) = 96 so you'll plug that into the equation 96 = C *(4)^2 Can you solve for C?

OpenStudy (anonymous):

what does the arrow between the 4 and the 2 mean?

OpenStudy (anonymous):

oh wait its an exponet

OpenStudy (anonymous):

so c=6

OpenStudy (anonymous):

Yes. So you have f(x) = 6x^2 that gets you f(2) f(2) = 6 * (2)^2

OpenStudy (anonymous):

ok I GOT IT thanks:):):) , do you mind helping me with another?

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

opps hold on

OpenStudy (anonymous):

I only have 10 minutes before I have to leave to my dance class

OpenStudy (anonymous):

sorry i got kicked out

OpenStudy (anonymous):

lol its okay

OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x^2, and g(x) = 144 when x = 6. answers: g(x)=24x^2 g(x)=6x^2 g(x)=4x^2 g(x)=1/4x^2

OpenStudy (anonymous):

sorry, I have to go, I wasn't able to finish solving the problem, I'm REALLY REALLY SORRY!

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

thanks anyway!!! ::))